# Performance of a coarsely pixelated LAPPD photosensor for the SoLID light gas Cherenkov detector

### v2

@@ Abstract
The SoLID spectrometer light gas Cherenkov counter requires photosensors that operate in a high luminosity and high background environment. In its reference design, a configuration of 9 or 16 tiled multi-anode photomultipliers (MaPMTs) covers the sensing area. To evaluate the use of a pixelated INCOM Large Area Picosecond Photodetector (LAPPD) as an alternative to cover the same area, we tested the performance of this photosensor in realistic SoLID running conditions in Hall C at the Thomas Jefferson National Accelerator Facility (TJNAF, a.k.a Jefferson Lab or JLab). The results of this test confirmed that the coarse-pixelated (2.5x2.5 cm pixel size) LAPPD is capable of handling the total projected signal and background rates of the three pillar experiments of SoLID. The photosensor detected Cherenkov signals with the capability of separating single-electron events from pair production events while rejecting background. Although the design was not aimed at Ring Imaging Cherenkovs, Cherenkov disk images were captured in two gas radiators. Through a direct comparison with a GEANT4 simulation, we confirmed the experimental performance of the LAPPD.
@@ 1 Introduction
The Solenoidal Large Intensity Device (SoLID) [1, 2] is a new spectrometer designed to exploit the full potential of the 12 GeV energy upgrade of the Continuous Electron Beam Accelerator Facility (CEBAF) at Jefferson Lab (JLab). The SoLID spectrometer will be operated at the full CEBAF luminosity of 10 cms [2] in the Parity Violation Deep Inelastic Scattering (PVDIS) experiment and at a reduced luminosity of 10 cms in the Semi-Inclusive Deep Inelastic Scattering (SIDIS) or Exclusive production near threshold, known as the SIDIS-J/psi experiments. In the SIDIS- SoLID spectrometer configuration, Cherenkov counters were designed for a trigger system and particle identification: a light gas Cherenkov (LGC) for e/ separation and a heavy gas Cherenkov (HGC) for /K and /p separation. Hamamatsu flat panel multi-anode photomultipliers (MaPMTs) [3] were selected as the baseline photosensors. The MaPMTs have 64 pixels with sensitivity down to single photon detection. Signals from pixels can be either read out individually (64 pixels) or as 16 quad-sum channels. However, a major drawback of this sensor is the loss of efficiency in a magnetic field [4], and thus magnetic shielding is unavoidable for it to be used in a magnetic field. An alternative sensor is the Microchannel plate photomultiplier (MCP-PMT), a photosensor with a proven strong magnetic field resiliency, and high granularity. Its design consists of a stacked pair of MCPs in a chevron configuration for secondary electron amplification. Due to the significantly reduced electron amplification path, the MCP-PMT shows an exceptional magnetic field tolerance [5], a great advantage to avoid heavy and cumbersome mu-metal magnetic shielding in large array configurations of MAPMTs in each counter sector of SoLID. Unfortunately, the high cost of these commercial MCP-PMTs prevents us from readily tiling them for large area coverage Cherenkov counters in the presence of strong magnetic fields.
@@
Over the past decade a new type of MCP-PMT, the large area picosecond photodetector (LAPPD) was developed by the LAPPD collaboration [5] and more recently commercialized by Incom, Inc [7, 8, 9]. The LAPPDs apply resistive and secondary emissive layers on a large area of glass capillary through atomic layer deposition (ALD), achieving low-cost functional MCPs, thus the final cost of the manufactured device is expected to be much lower per active area than that of traditional MCP-PMTs. Meanwhile, the performance of the LAPPD is comparable to that of the standard commercial MCP-PMTs, including timing resolution, position resolution, and dark count. The expected low cost and high performance make the LAPPD an ideal photosensor candidate for Cherenkov counters.
@@
To explore the potential of using a LAPPD as the photosensor for the SoLID Cherenkov counter and evaluate its performance in a high-rate background environment, we performed an earlier opportunistic but limited test [10] of a strip-line read-out Gen-I LAPPD at Jefferson Lab. Although the event separation was not very clean due to a low quantum efficiency and non-pixelated readout the results showed a promising performance of the GEN-I LAPPD in a harsh running environment of Hall C at Jefferson Lab. With the availability of Gen-II (capacitive coupling) pixelated LAPPD, a more comprehensive test was designed and performed to validate it as a viable photosensor for the SoLID gas Cherenkov counters as part of a pre-R&D. In this paper, we describe the details of this comprehensive test, with the experimental setup, the Gen-II LAPPD geometry and data taking procedure. Results will be presented and discussed as we evaluate the potential of a coarse-pixelated LAPPD as a photosensor candidate for the SoLID gas Cherenkov counters.
@@ 2 Telescopic Cherenkov device design
A Telescopic Cherenkov Device (TCD) LAPPD was designed and constructed, as shown in Figure 1. Compared to Ref. [10] the present design to allow the use of an array of 4x4 MaPMTs or a 20x20 cm LAPPD. It consists of the primary Cherenkov tank, the LAPPD electronics housing box, the aluminum-coated reflective mirror, aluminum foil end caps, the gas supply system, the power supply system, and readout electronics.
The primary Cherenkov tank was assembled from 14" diameter PVC pipes, and machine cut to the required dimensions. Inside the tank, felt flocking was attached to minimize secondary reflections. Additional apertures were added to allow valves and gas flow, sensors, and an LED test light source. A special housing box was designed to accommodate the coarse-pixelated LAPPD and its electronic cables. The housing box was designed to maintain structural and hermetic integrity while allowing detachable access plates in order to reach the enclosed electronics. To secure the removable LAPPD electronics housing box, large circular rubber gaskets were applied. Aluminum windows and PVC window frames were pressure tested to ensure the system was airtight. Flat elliptical mirrors were constructed from carbon-fiber bases and Lexan reflective film. The primary mirror was mounted inside the tank and aligned to reflect light incident down the central ray of the tank and reflected by the mirror onto the center of the photon detector array. The LAPPD housing box was surrounded with lead bricks during beamtime to protect the photosensor from radiation damage. A 370 nm light-emitting device (LED) was installed near the entrance window for calibration and checking purposes. A pressure regulation system was assembled, it included a gauge pressure manometer, two solenoid valves, a desktop PID controller with a laptop interface, and the required cabling. Operating at atmospheric pressure helped simplify the mechanical design and minimized the window thicknesses. Carbon dioxide (CO) and octafluorocyclobutane (CF) were used as the radiator medium in this system to evaluate the LAPPD performance for both light gas Cherenkov and heavy gas Cherenkov (see next section). In order to ensure high gas purity and before stabilizing it at atmospheric pressure, the telescopic Cherenkov device was flushed in several cycles each time a new gas was used.
@@ 3 Large area picosecond photodetector (LAPPD)
The tested Gen-II LAPPD features a 64-pixelated readout with a pixel size of 25 mm x 25 mm through capacitive coupling. The entrance window of this LAPPD is regular B33 glass which absorbs photons below 270 nm. To enhance the Cherenkov photons detection efficiency a lower wavelengths, a thin layer of p-terphenyl wavelength shifter coating was applied [11] to convert ultraviolet photons into visible photons for detection. Fig.1(a) shows the exploded-view drawing of LAPPD and an image of the tested Gen-II coarse-pixelated LAPPD installed in the telescopic Cherenkov housing box. The detailed parameters and performance on the bench of this particular LAPPD are listed in Table 1.
@@
| Device size | 20 20 cm |
| Active area | 20 20 cm |
@@ 4 Experimental setup
The telescopic Cherenkov device was installed in hall C at Jefferson Lab, along with the Super High Momentum Spectrometer (SHMS). The experiment was performed parasitically with the ongoing Hall C fixed-target experiments, the setup location was chosen for minimal interference with the running experiment as well as to have a moderate rate background environment. The Cherenkov was positioned to face the entrance window of the liquid hydrogen target to allow the scattered electrons from the target to freely pass through the entrance window after crossing the target and the target scattering chamber aluminum walls.
@@
The experiment used an 8.1 GeV electron beam, impinging a liquid H target. While the scattered electrons or produced particles pass through the tank, Cherenkov light is produced along the particle path if its velocity fulfills the Cherenkov light emission condition. The Cherenkov light is then reflected at 90 by the mirror at the section, and collected by the LAPPD. The coarse-pixelated LAPPD was accommodated in the housing box, and surrounded by lead blocks to prevent radiation damage. Scintillator paddles and Shashlik electromagnetic calorimeter blocks [12] were placed in front of the entrance window and behind the exit window respectively to generate coincident external triggers. All triggers and signals from the scintillator paddles, the calorimeter, and LAPPD were fanned out and sent to the F250 flash Analog-to-Digital Converters (FADC250) [13] for data acquisition.
@@
The current SoLID detector design includes two Cherenkov counter sub-systems: the light gas Cherenkov (LGC) and the heavy gas Cherenkov (HGC). The light gas Cherenkov applies CO at 1 atm (index of refraction n = 1.00045) as a gas radiator, it will be mainly used to separate electrons from pions giving a very low momentum threshold for electrons (0.017 GeV) and a medium momentum threshold for pions (4.75 GeV). The heavy gas Cherenkov applies CF at 1.7 atm (index of refraction n = xxx) as gas radiator, it will provide a clear distinction between charged pion and kaon Cherenkov signals from a momentum range of 2.5 GeV/c to 7.5 GeV/c for the SoLID SIDIS program at forward angles from 8 to 15. To evaluate the performance of pixelated LAPPD in both the light gas and heavy gas Cherenkov sub-systems, we experimented with CO and CF as the gas radiators, respectively.
@@
The Cherenkov counter was installed in Hall C at Jefferson Lab, as shown in Figure 3. It was mounted at beamline level pointing towards the target cell and positioned 4.8 m away from the target scattering chamber. Its entrance window was facing the target at an angle of 68 with respect to the beam direction. The distance and angle were constrained by space and safety considerations due to the ongoing experiment in the hall, resulting in a moderate background-rate environment. Lead blocks shielded the photosensors from direct radiation damage.
@@ 5 Simulation
In order to understand the experimental test results and later optimize the SoLID LGC design, we conducted a thorough simulation of the entire Cherenkov process. A Geant4 simulation [14] of the TCD, including the Cherenkov counter, the EM calorimeter (EC), and scintillator (SC) detectors, was developed similar to the SoLID simulation. The setup is shown in Figure 4. The measured LAPPD quantum efficiency was used and the Cherenkov optical physics cut-off was chosen to be 200 nm. We studied the resulting number of photo-electrons in the Cherenkov detector from two main sources, namely high energy scattered electrons and /0 production with subsequent decay photons conversion into electron-positron pairs from the polarized He gas target. Simulation results are presented together with the experimental test results for comparison in the following sections.
The TCD setup in the Geant4 simulation.
@@ 6 Results and Discussion
The LAPPD has a signal readout scheme of 8 × 8 array of 25 mm × 25 mm pixels. Signals from all 64 channels were recorded during the experiment, and the data was analyzed offline. In a Cherenkov signal event, the generated photons will be distributed as a light disk at the LAPPD photosensor. Multiple pixels should collect Cherenkov photons and fire simultaneously. Figure 5 shows the signal sum for different numbers of pixels (N) subgroups. The signals can be categorized into three groups: i) the random coincidence of single photo-electron signals with N = 1, 2, 3, demonstrating a significant suppression when requiring multiple pixels (channels) to fire within the same time window. ii) Cherenkov signals with N = 9, 10, 11. The signals follow a Gaussian distribution centered around the number of photoelectrons (NPE) of 17, corresponding to central hits. iii) Gaussian-like signals with N = 19, 20, 21, approximately double the NPE of the typical Cherenkov events. These high-NPE signals are possibly the coincidence sums of simultaneous detection of an electron and positron from a pair production process.
The LAPPD test includes different types of radiator gas: CO, a mixture of 80% CF and 20% CO, and CF. The normalized distributions of the signal sum from these types of gases are shown in Figure 6. All three distributions show an identical peak below 100 ADC channel, indicating a stable random coincidence background from the high rate environment during the beam test. The heavy gas (CF) shows a more prominent bump from 250 to 1000 ADC channels, indicating more photons detected from typical Cherenkov events.
An important feature of the Cherenkov counter is the detection of the photon disk with photosensors. The disk size is related to the particle’s momentum, the refraction index of the radiator, and the photon’s path length to the photosensors. With the light gas radiator (CO), a typical Cherenkov photon ring would have 9 to 10 fired pixels. In heavy gas (CF), the Cherenkov light cone contains more photons and has a larger opening angle, resulting in a greater disk radius on the detection plane with more fired pixels. According to their refractive index (n = 1.0005 and n = 1.0014), the Cherenkov disk from CF is about 70% larger in radius than the disk from CO. Thus, the corresponding coverage of a typical Cherenkov signal is 15 to 16 pixels in the heavy gas. The top panel of Figure 7 shows the Cherenkov rings we observed with the LAPPD and different gas radiators from the large scattering angle beam test. The event samples show a small shift of the signal center to the right of the geometrical center and a slightly wider distribution, compared to the simulation shown in the bottom panel of Figure 7. Without a tracking detector, we cannot precisely select the events that enter the telescope Cherenkov detector right on its geometrical center, so shifts from event to event are expected. Most importantly, the size and shape of the Chereknov rings observed from the data are consistent with those in the simulation.
Given the expected light disk sizes for typical Cherenkov events in different gas radiators, we sum all the signals with the requirement of 9 to 10 fired pixels in CO and 15 to 16 in CF. The signal sums with the LAPPD sensor show an NPE = 12 in CO and NPE = 21 in CF. Our simulation can also well describe the difference 20 between the two gas radiators. The simulation yields NPE = 9 in CO and NPE = 23 in C4F8, showing a good agreement with the data. The results are presented in Figure 8.

### v4

# Performance of a coarsely pixelated LAPPD photosensor for the SoLID gas Cherenkov detectors

@@ Abstract
The SoLID spectrometer’s gas Cherenkov counters require photosensors that operate in a high luminosity and high background environment. The reference design features arrays of 9 or 16 tiled multi-anode photomultipliers (MaPMTs), distributed across 32 sectors, to serve the light-gas and heavy-gas Cherenkov counters, respectively. To assess the viability of a pixelated INCOM Large Area Picosecond Photodetector (LAPPDTM) as an alternative photosensor to replace MaPMT arrays in either detector, we evaluated its performance under realistic SoLID running conditions in Hall C at the Thomas Jefferson National Accelerator Facility (Jefferson Lab).
The results of this test confirmed that the coarse-pixelated (2.52.5 cm2 pixel size) LAPPD is capable of handling the total projected signal and background rates of the three pillar SoLID experiments. The tested photosensor detected Cherenkov signals with the capability of separating single-electron events from pair production events while rejecting background. Although the design was not aimed at ring-imaging Cherenkov detectors, Cherenkov disk images were captured in two different gas radiators. Through a direct comparison with a GEANT4 simulation, we confirmed the experimental performance of the LAPPD.
@@ 1 Introduction
The Solenoidal Large Intensity Device (SoLID) [1, 2, 3] is a new spectrometer designed to exploit the full potential of the 12 GeV energy upgrade of the Continuous Electron Beam Accelerator Facility (CEBAF) at Jefferson Lab. The SoLID spectrometer will be operated at the full CEBAF luminosity of 1039 cm-2s-1 [3] in the Parity Violation Deep Inelastic Scattering (PVDIS) experiment and at a still high but reduced luminosity of 1037 cm-2s-1 in the Semi-Inclusive Deep Inelastic Scattering (SIDIS) and Exclusive production near threshold experiments (the SIDIS- experiments).
@@
The SIDIS- spectrometer design uses two Cherenkov counters for particle identification: a light gas Cherenkov (LGC) for e/ separation and a heavy gas Cherenkov (HGC) for and separation. Hamamatsu flat panel multi-anode photomultipliers (MaPMTs) [4] were selected as the baseline photosensors for these Cherenkov detectors. The MaPMTs have 64 pixels with a sensitivity down to signals from single photons. The signals from the pixels can be read out either individually (64 pixels) or in various summing schemes, e.g. summing 16 pixels into four quadrants. However, a major drawback of these sensors is their loss of efficiency in a magnetic field [5, 6], requiring the use of magnetic shielding over the full photosensing area for their use in a magnetic field. An alternative sensor is the microchannel plate photomultiplier (MCP-PMT), a photosensor with a proven strong magnetic field resiliency, and high granularity. Its design consists of a stacked pair of MCPs in a chevron configuration for secondary electron amplification. Due to the significantly reduced electron amplification path, MCP-PMTs show an exceptional magnetic field tolerance [7], a great advantage compared to the heavy and cumbersome mu-metal magnetic shielding needed for large array configurations of MAPMTs in the SoLID Cherenkov detectors. Unfortunately, the prohibitive cost of these commercial MCP-PMTs prevents their use in large photosensing arrays.
@@
Over the past decade a new type of MCP-PMT, the large area picosecond photodetector (LAPPD) was developed by the LAPPD collaboration [8] and more recently commercialized by Incom, Inc [9, 10, 11]. The LAPPDs apply resistive and secondary emissive layers on a large area of glass capillary through atomic layer deposition (ALD), achieving low-cost functional MCPs. As a result the final cost of a manufactured device is expected to be much lower per active photosensing area compared to traditional MCP-PMTs, meanwhile, the performance of the LAPPD is similar to that of the standard commercial MCP-PMTs, including timing resolution, position resolution, and dark count. The expected low cost and high performance make the LAPPD an ideal photosensor candidate for Cherenkov counters.
@@
To explore the potential of using LAPPDs as the photosensors for the SoLID Cherenkov counters and to evaluate its performance in a high-rate background environment, we performed a first opportunistic but limited test [12] of a strip-line read-out Gen-I LAPPD at Jefferson Lab. Although the event separation between single-electron and pair-production events was poor due to a low quantum efficiency and non-pixelated readout, the results showed a promising performance of the GEN-I LAPPD in a harsh running environment of Hall C at Jefferson Lab. With the availability of Gen-II (capacitive coupling) pixelated LAPPD, a more comprehensive test was designed and performed to validate it as a viable photosensor for the SoLID gas Cherenkov counters as part of a pre-R&D. In this paper, we describe the details of this comprehensive test, including the experimental setup, the Gen-II LAPPD geometry and the data taking procedure. Results are presented and discussed as we evaluate the potential of a coarse-pixelated LAPPD as a possible photosensor candidate for the SoLID gas Cherenkov counters.
@@ 2 Telescopic Cherenkov device design
A Telescopic Cherenkov Device (TCD) was designed and constructed, as shown in Figure 1. Compared to Ref. [12] the present design allows the use of an array of 44 MaPMTs or a 2020 cm2 LAPPD, the size of the sensor array in the SoLID HGC. It consists of a primary Cherenkov tank, a LAPPD electronics housing box, an aluminum-coated reflective mirror, aluminum foil end caps, a gas supply system, a power supply system, and readout electronics.
The primary Cherenkov tank was assembled from 14" diameter PVC pipes, and machine cut to the required dimensions. Inside the tank, felt flocking was attached to minimize secondary reflections. Additional apertures were added to allow valves and gas flow, sensors, and an LED test light source. A special housing box was designed to accommodate the coarse-pixelated LAPPD and its electronic cables. The housing box was designed to maintain structural and hermetic integrity while allowing detachable access plates in order to reach the enclosed electronics. To secure the removable LAPPD electronics housing box, large square rubber gaskets were applied. Aluminum windows and PVC window frames were pressure tested to ensure the system was airtight. Flat elliptical mirrors were constructed from carbon-fiber bases and Lexan reflective film. The primary mirror was mounted inside the tank and aligned to reflect light incident down the central ray of the tank and reflected by the mirror onto the center of the photon detector array. The LAPPD housing box was shielded by lead bricks during data taking to protect the photosensor from radiation damage. A 370 nm light-emitting diode (LED) was installed near the entrance window for calibration and monitoring purposes. The pressure regulation system included a gauge pressure manometer, two solenoid valves, a desktop PID controller with a laptop interface, and the required cabling. Operating slightly above atmospheric pressure helped simplify the mechanical design and minimize the entrance window thickness. Carbon dioxide (CO2) and octafluorocyclobutane (C4F8) were used as radiators media to evaluate the LAPPD performance for both the light gas (CO2) and heavy gas (C4F8) Cherenkov (see next section). In order to ensure high gas purity and before stabilizing it at atmospheric pressure, the telescopic Cherenkov device was flushed with 1 to 2 full detector equivalent volumes of the new gas each time the radiator was changed.
@@ 3 Large area picosecond photodetector (LAPPD)
The tested Gen-II LAPPD features a 64-pixelated readout with a pixel size of 25 mm 25 mm through capacitive coupling. The entrance window of this LAPPD is regular B33 glass which absorbs photons below 270 nm. To enhance the Cherenkov photons detection efficiency at lower wavelengths, a thin layer of p-terphenyl wavelength shifter coating was applied [13] in order to convert ultraviolet photons into visible photons for detection. Fig.1 shows the exploded-view drawing of a LAPPD and an image of the tested coarse-pixelated LAPPD installed in the telescopic Cherenkov housing box, as well as its quantum efficiency map at 365 nm. The detailed parameters and performance on the bench of this particular LAPPD are listed in Table 1.
@@
| Device size | 20 20 cm2 |
| Active area | 20 20 cm2 |
@@ 4 Experimental setup
The telescopic Cherenkov device was installed in Hall C at Jefferson Lab, along with the Super High Momentum Spectrometer (SHMS). The experiment was performed parasitically with the ongoing Hall C fixed-target experiments, the setup location was chosen for minimal interference with the running experiment as well as to have a moderate rate background environment. The Cherenkov was positioned to face the liquid hydrogen target to allow the scattered electrons from the target to freely pass through the Cherenkov entrance window after crossing the target side walls and the target scattering chamber aluminum walls.
@@
The experiment used a 10.4 GeV electron beam, impinging on a polarized gaseous 3He target. While the scattered electrons or produced particles passed through the tank, Cherenkov light is produced along the particle path if its velocity fulfills the Cherenkov light emission condition. The Cherenkov light is then reflected at 90∘ by the mirror at the section, and collected by the LAPPD. The coarse-pixelated LAPPD was accommodated in the housing box, and surrounded by lead blocks to prevent radiation damage. Scintillator paddles and Shashlik electromagnetic calorimeter blocks [14] were placed in front of the entrance window and behind the exit window respectively to generate coincident external triggers. All triggers and signals from the scintillator paddles, the calorimeter, and LAPPD were fanned out and sent to F250 flash Analog-to-Digital Converters (FADC250) [15] for data acquisition.
@@
The current SoLID detector design includes two Cherenkov counters: the LGC and the HGC. The LGC uses CO2 at 1 atm (index of refraction n = 1.00045) as a gas radiator, and will be mainly used to discriminate electrons (17 MeV Cherenkov light threshold) from pions (4.75 GeV Cherenkov light threshold). The HGC, critical for the SoLID SIDIS program at forward angles from 8∘ to 15∘, uses C4F8 at 1.7 atm (index of refraction n = 1.0014) as a gas radiator and provides a clear separation between charged pions and kaons Cherenkov signals in a momentum range between 2.5 GeV/c and 7.5 GeV/c . To evaluate the performance of pixelated LAPPD in both the light gas and heavy gas Cherenkov, we experimented with CO2 and C4F8 as the gas radiators, respectively. We conducted these tests in addition to rate dependence studies with the baseline Ma-PMT arrays.
@@
The Cherenkov counter was installed in Hall C at Jefferson Lab, as shown in Figure 3. It was mounted at beamline level pointing towards the target cell and positioned 5.4 m away. Its entrance window was facing the target at an angle of 75∘ with respect to the beam direction. The distance and angle were constrained by space and safety considerations due to the ongoing experiment in the hall, resulting in a moderate background-rate environment. Lead blocks shielded the photosensors from direct radiation damage.
We exposed the pixelated LAPPD to rates of up to 21 kHz per pixel, consistent with the projected rates for the SoLID SIDIS- conditions. During the same test period, we evaluated the MaPMT arrays at smaller angles under rate conditions with 0.25 photoelectron rates exceeding 3.4 MHz per quadrant, surpassing the projected single photon background rate of 7.1 MHz per MaPMT for the SoLID-PVDIS configuration. Previous studies on strip-line LAPPDs [12] have demonstrated their robustness in high-rate environments, suggesting that LAPPDs, like MaPMTs, are suitable for the harsh PVDIS conditions. Consequently, LAPPDs present a viable alternative to MaPMTs for the SoLID gas Cherenkov detectors.
@@ 5 Simulation
In order to understand the experimental test results and later optimize the SoLID LGC design, we conducted a thorough simulation of the entire Cherenkov process. A Geant4 simulation [16] of the TCD, including the Cherenkov counter, the EM calorimeter (EC), and scintillator (SC) detectors, was developed similar to the SoLID simulation. The TCD setup in our simulation is shown in Figure 4. The measured LAPPD quantum efficiency was used and a conservative Cherenkov optical physics cut-off was chosen to be 200 nm.
We studied the resulting number of photo-electrons in the Cherenkov detector from two main sources, namely high energy scattered electrons and ’s production with subsequent decay photons conversion into electron-positron pairs from the polarized 3He gas target. Simulation results are presented together with the experimental test results for comparison in the following section.
@@ 6 Results and Discussion
The LAPPD utilizes a readout scheme featuring an array of 88 pixels, each reading a 25 mm 25 mm area. Throughout the beam tests, various radiator gas mixtures were employed, including CO2, an 80% C4F8 / 20% CO2 blend, and pure C4F8. Signals from all 64 pixels were recorded, and the analysis was conducted offline. Figure 5 illustrates the normalized distributions of the aggregated signal across all 64 pixels. A coincidence timing cut of ns has been applied to the relative timing between pixels and the triggering calorimeter block. All three distributions exhibit a consistent peak below 100 ADC channels, indicating a stable background of random coincidences arising from the high-rate environment. Notably, signals in the heavier gas, C4F8, show a prominent signal amplitude bump between ADC channel 500 and 750, most likely a result of the Cherenkov radiation in C4F8.
The single photoelectron (SPE) responses were analyzed using data collected outside the timing coincidence window. This analysis revealed that the SPE peak consistently ranged between ADC channels 16 and 20 across all pixels. Given that the signal amplitude for the majority of our data does not exceed 2 SPE, we assumed a linear response. Consequently, all signals were normalized to the SPE amplitude to extract the number of photoelectrons (NPE).
As depicted in Figure 5, the data are overwhelmed by the random coincidences background at low ADC channels, an expected phenomenon caused by the high-rate environment experienced during the test. Thus, distinguishing Cherenkov signals from this background presents a significant challenge, particularly for the light gas radiator (CO2) because of its relatively low light yield. To effectively select the Cherenkov signals, it is imperative to leverage the granularity afforded by the pixelized readout. Typically, photons from a Cherenkov event are distributed in an annular pattern on the photosensor plane, illuminating multiple pixels simultaneously. The exact number of activated pixels varies depending on the photon annulus’s location and size, which are affected by the incident particle’s angle and momentum, the refractive index of the radiator, and the photon’s path length to the photosensors.
The simulation results for Cherenkov photon distributions are shown in Figure 6. For CO2 gas (with a refractive index = 1.00045), the photons annulus covers 12 pixels when they are centered at one pixel, activating roughly 8 pixels given the LAPPD’s quantum efficiency. In the case of C4F8 gas radiator ( = 1.0014), the Cherenkov light cone contains a higher photons count and exhibits a larger opening angle, resulting in a coverage of 32 pixels and approximately 15 activated pixels if the trajectories of the incident electrons coincide with the TCD’s centerline. Here a pixel is considered "activated" if its signal exceeds 0.5 NPE, which is an arbitrary choice to reject dark current signals. Our results are stable by varying this threshold from 0.4 to 0.6 NPE.
The shape of the Cherenkov photon annulus is also studied with beam test data. To select clean samples of Cherenkov events, square annulus cuts were applied to the data. The annulus is one pixel wide and its center is defined as the bottom right corner of the inner square. The cut can be applied on the center pixel (at ith row and jth column of the array) and its surrounding pixels with three parameters: the size of the inner square (), the number of inactivated pixels inside the inner square (), and the number of activated pixels within the annulus (). Table 2 lists two sets of parameters chosen for different gas radiators. The photon distribution with the annulus cuts on for both CO2 and C4F8 data are shown in Figure 7. The observed size and shape of the photon annulus in the data agree well with the predictions from our simulations.
| Cut | Gas Radiator | Center | |||
| Annulus1 | CO2 | (out of 1) | (out of 8) | ||
| Annulus2 | C4F8 | (out of 4) | (out of 12) |
The annulus cuts are then applied across all the possible centers (see Table 2) on the photosensor to improve the statistics of the Cherenkov event samples. Figure 8 shows the 2D distribution of the selected events over the NPE and the number of activated pixels (Npix). For both CO2 and C4F8 data, a clear peak is revealed after the cut, showing that a typical Cherenkov event would activate up to 11 pixels in CO2 and 17 pixels in C4F8 in the TCD.
The NPE distributions from the data and the simulation are shown in Figure 9. The same annulus cuts are used for the beam test data. In addition, we also require the number of activated pixels outside the annulus shape to be less than that on the annulus, which significantly suppresses the large NPE tail on the distribution plot. This distribution characterizes the LAPPD responses to the Cherenkov photons in TCD, giving for CO2 and for C4F8. The simulation also describes the difference between the two gas radiators, yielding an NPE of 9 for CO2 and 23 for C4F8. The beam test data and simulation are consistent with each other on the NPE distribution of the Cherenkov signals.